OpenStudy (anonymous):

Suppose a and b are real numbers, not both 0. Find real numbers c and d such that 1/(a+bi)=c+di

OpenStudy (bobo-i-bo):

Let me start you off: \[\frac{ 1}{ a+bi } = \frac{ 1 }{ a+bi }\frac{ a-bi }{ a-bi }=?\]